Select financial information follows. Calculate Cash Flow from Assets, Hint: In the process, you need to calculate the amount of taxes that would appear on the income statement based on the information available to you. Round your answer to 2 decimal places (example if your answer is 546.345 you should enter 546.35). Margin of error for correct responses:+/-$.05 Note: If your final answer is negative, be sure to enter a negative sign with your respons from from from Income Balance Balance tatement, Sheet Sheet Year ending 2/31/2031 12/81/2092 12/31/202 Depr 192 current assets 151 178 Expense net fixedR assets640EBIT 980 nterest 101 current liabilities 147 expense Taxes (at? 30%)
Answer #1

Answer:

Working capital in 2031= current assets - current liabilities =151-147=$4

Working capital in 2032= current assets - current liabilities =178-126=$52

Change in working capital =52-4=$48

Investment in Fixed asset =Net fixed asset in 2032-(net Fixed asset in 2031-depreciation)=571-(640-192)=$123

So cash flow

Depreciation EXP D 192
EBIT 980
Interest expenses I 101
EBT=EBIT-I 879
Taxes @ 30%  T=30%*EBT 263.7
PAT=EBT-T 615.3
Change in working capital W 48
Change in capital Exp C 123
Net Cash flow=PAT+D-W-C 636.3